OpenStudy (anonymous):

A student opens a mathematics book to two facing pages. The product of the pages is 420. Find the page numbers.

OpenStudy (anonymous):

let the first page no be =x so the second page no will be =x+1 so (x+1)x=420 solve this equation for x, you will find the answer

OpenStudy (anonymous):

Ok, ravi623. I come up with x^2=420-x? But what do I do after this?, because x is negative do I need to make it positive first? -1(x-420) instead?

OpenStudy (anonymous):

factorize the the 420=21X20 (x+1)x=420 becomes x^2+x-420=0, this becomes a quadratic equation so x^2+21x-20x-420=0 x(x+21)-20(x+21)=0 (x-20)(x+21)=0 so x=20, x=-21 but no of pages can not be negative ,so x=20 is the right answer so first page no is 20 and other pages no is 20+1=21

OpenStudy (anonymous):

Thank you, I understand